OpenStudy (anonymous):

I am trying to figure out how to solve this equation using (pemdas) (4x2)-(1x9). I keep getting 8-9= which can't be solved right?

OpenStudy (anonymous):

8-9=-1 isnt it?

OpenStudy (anonymous):

Ok well can you guide me in the right direction?

OpenStudy (anonymous):

i can try:)

OpenStudy (anonymous):

whats prob r u facing?

OpenStudy (anonymous):

I am so terrible at math and I feel like I need to multiply what is grouped first and then subtract.

OpenStudy (anonymous):

yes u r right, first solve the expression in brackets, then subtract

OpenStudy (anonymous):

ok so would the answer be 8

OpenStudy (anonymous):

8-0=8

OpenStudy (anonymous):

8-0=0 for sure but here u have 8-9=-1

OpenStudy (anonymous):

I didn't pay attention to 1x9=0 I was looking at it the other way around.

OpenStudy (anonymous):

1*9=9

OpenStudy (anonymous):

yes. thanks!

OpenStudy (shadowfiend):

Remember to give people a medal when they're helpful :)

OpenStudy (anonymous):

Thanks i'm new to this!

OpenStudy (shadowfiend):

No worries :) Hahaha. I don't know that *I* deserved a medal there, but I appreciate it!

OpenStudy (anonymous):

No prob:)